Solve equation using elmenation.
2x + 3y = 0
3x + y = 7

Answer by TimothyLamb(4379)

L1: 2x + 3y = 0
L2: 3x + y = 7
L2: -9x - 3y = -21
L1 + L2:
-7x = -21
x = 3
y = -2
Answer: (3, -2)
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php